OpenStudy (anonymous):

Evaluate the integral 75Cubed root of x dx

zepdrix (zepdrix):

Remember how to write roots as rational expressions? Example:\[\large \sqrt x \qquad = \qquad x^{1/2}\] We've changed it to an exponent. The denominator represents the degree of the root. \[\large \int\limits 75\sqrt[3]{x}\;dx \qquad = \qquad 75\int\limits x^{1/3}\;dx\] Understand that part?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

From here, just do as you've done in previous problems. Apply the `Power Rule for Integration`.\[\large 75\int\limits\limits x^{1/3}\;dx \qquad = \qquad 75 \frac{x^{(1/3+1)}}{1/3+1}+C\]

OpenStudy (anonymous):

ok thank you